A).a circular steel wire 2.00m long must stretch no more than 0.25cm when a tensile force of 400N is applied to each end of the wire .What minimum diameter is required for the wire ?
1
Expert's answer
2020-06-23T12:59:44-0400

Determine the strain:


"\epsilon=\frac{\Delta l}{l}."

Determine the stress:


"\sigma=E\epsilon=E\frac{\Delta l}{l}."

Determine the cross-sectional area:


"A=\frac{F}{\sigma}=\frac{Fl}{E\Delta l}."

Determine the diameter:


"d=2\sqrt{\frac{A}{\pi}}=2\sqrt{\frac{Fl}{E\Delta l\pi}}."

Substitute the values (take Young's modulus for steel E=200 GPa):


"A=2\sqrt{\frac{400\cdot2}{200\cdot10^9\cdot0.0025\pi}}=1.42\cdot10^{-3}\text{ m}."